Liam thinks of a number.
He multiplies the number by 5 and then subtracts 60 from the result .
His answer equals the number he started with .
What was the number Liam started with ?

Answer by nerdybill(6958)

Liam thinks of a number.
He multiplies the number by 5 and then subtracts 60 from the result .
His answer equals the number he started with .
.
Let x = number Liam started with
then
5x-60 = x
5x = x+60
4x = 60
x = 60/4
x = 15
.
What was the number Liam started with ? 15